An incident ray strikes a glass (n=1.52) with an angle of 380 to the surface of the glass.

A. Illustrate with label angle of incidence, angle of reflected ray and angle of refraction.

B. What is the angle of reflected ray?


1
Expert's answer
2022-05-19T05:42:33-0400

A.



B.


"n_i\sin\theta_i=n_r\sin\theta_r"

"\theta_r=\sin^{-1}\frac{1.00*\sin38^\circ}{1.52}=24^\circ"